# DELTA: Dense Efficient Long-range 3D Tracking for any video

### v1

@@ Abstract
Figure 1: DELTA is a dense 3D tracking approach that (a) tracks every pixel from a monocular video, (b) provides consistent trajectories in 3D space, and (c) achieves state-of-the-art accuracy on 3D tracking benchmarks while being significantly faster than previous approaches in the dense setting.
@@ 2 Related Work
Optical Flow estimates motion by providing dense pixel-wise correspondences between consecutive frames. Early variational (Mémin & Pérez, 1998; Horn & Schunck, 1980; Brox et al., 2004) struggled with robustness in complex scenes with rapid motion, occlusions, and large displacements. The introduction of CNN-based methods (Ilg et al., 2017; Ranjan & Black, 2017; Xu et al., 2017; Sun et al., 2018) improved motion estimation between adjacent frames. RAFT (Teed & Deng, 2020) marked a breakthrough by leveraging 4D correlation volumes for all pairs of pixels. Follow-up works advanced this by incorporating transformers for tokenizing 4D correlation volumes (Huang et al., 2022), adopting global motion feature aggregation to improve prediction in occluded regions (Jiang et al., 2021), and framing optical flow as a matching problem with correlation softmax operations (Xu et al., 2022). While some efforts propose to apply optical flow to long-term sequences with multi-frame optical flow (Teed & Deng, 2020; Godet et al., 2021; Shi et al., 2023) or integration of point-tracking techniques (Le Moing et al., 2024; Cho et al., 2024a), they often suffer from drifting and occlusion challenges, limiting their reliability for long-term tracking.
@@
Among these methods, we are the first feed-forward approach that performs dense 3D tracking efficiently from a long-term and in-the-wild video.  Table 1 provides a brief comparison of existing motion estimation methods alongside our approach.
@@ 3.1 Preliminary: point tracking
Limitations in spatial attention. The primary bottleneck in prior point tracking methods for dense tracking lies in spatial attention, which is sub-optimal due to the following limitations.
Computational complexity. As shown in Fig. 3, naively applying spatial self-attention across all tokens in a frame results in a complexity of $O(TN^{2})$, where $N$ is the number of tracks, making it impractical for dense tracking. To reduce complexity, CoTracker introduces virtual track tokens which are conceptually similar to learnable tokens introduced by DETR (Carion et al., 2020). It performs cross-attentions back-and-forth between trajectory tokens and a small number of virtual tracks, and self-attention is only applied within the virtual tracks. This reduces the complexity to $O(TKN+TK^{2}+KT^{2})$, assuming that the number of virtual tracks $K\ll N$. However, this reduction is still not enough for end-to-end tracking of every pixel in a high-resolution video. Although in practice, pixels could be divided into disjoint groups and perform attention within each group separately, this strategy is sub-optimal due to the interaction between tokens from different groups are ignored.
@@
Method overview. To further reduce computation cost without sacrificing accuracy, we adopt a strategy commonly used in optical flow (Teed & Deng, 2020; Huang et al., 2022): performing complex computations at the reduced spatial resolution, followed by lighter layers for upsampling. As shown in Fig. 2, we first run dense tracking on 1/$r^{2}$ of the original resolution, reducing the computation cost by 1/$r^{2}$. Then the reduced-resolution tracks are upsampled to full spatial resolution.
In Sec 3.2, we first discuss our design for reduced-resolution tracking with a new spatial attention architecture, which maintains linear complexity w.r.t. number of tracks while providing finer spatial granularity compared to CoTracker. Next in Sec 3.3 we introduce a new transformer-based upsampler that effectively predicts high-res tracking. Finally in Sec 3.4, we delve into details of depth representation that is crucial for accurate tracking in 3D.
@@ 3.2 Joint global and local spatial attention for efficient dense tracking
The most efficient spatial attention architecture from previous works, i.e. CoTracker (Karaev et al., 2023), has linear complexity w.r.t. number of tracks, enabling inference at reduced resolution on a single GPU, but training is still too costly. Moreover, the limited number of virtual tracks restricts the capture of fine-grained spatial relations, which hinders accuracy, particularly in dense tracking.
To address these limitations, we propose the following designs (illusted in Fig. 3) for an efficient spatial attention architecture that preserves fine-grained spatial relations while allowing efficient end-to-end training.
Global attention with sparse anchor tracks. We retain a small number of virtual tracks to represent the coarse global structure, cross-attending them only to a sparse set of anchor tracks, whose starting positions are uniformly sampled from the first frame. This approach reduces the computation cost of spatial attention by nearly half, assuming a small number of anchor tracks. As shown in figure 3 and Table 6(b), this reduction in computation does not noticeably impact tracking accuracy.
More importantly, this approach enables efficient training of dense tracking through a patch-by-patch strategy, without causing a training-testing time mismatch. During training, the virtual tracks first cross-attend a small number of anchor tracks and then cross-attend only a local random patch of tracks, rather than the entire image, to reduce computational costs. Since the virtual tracks are always computed from the same global anchor tracks, regardless of how the patch is cropped, this strategy maintains consistency. This is in contrast to CoTracker, where virtual tracks vary across different groups of tracks.
Patch-wise dense local attention. To capture fine-grained representations of local relations, we apply self-attention within small spatial patches before cross-attending to virtual tracks. Since self-attention is applied only within small patches, this approach adds a marginal complexity of $O(TNL)$ during inference, where $L\ll N$ is the number of tracks per patch. Experiments (Table 6(b)) show that incorporating dense local attention significantly improves dense tracking accuracy.
@@ 3.3 High-resolution track upsampler
Given the tracks extracted at the reduced spatial resolution $\frac{H}{r}\times\frac{W}{r}$ from our previous module, our next step is to upsample them to full resolution $H\times W$ in a single feedforward pass. In the context of optical flow, a common upsampling approach expresses the flow for each fine-resolution pixel as a convex combination of its nearest neighbor flows estimated in a coarse resolution (Teed & Deng, 2020). The weights for the combination are learned via a convnet. On the contrary, we instead introduce an attention-based mechanism for upsampling. We demonstrate the efficacy in our experiments (see figure 4 and Table 6(c)).
@@
Starting from the first input frame of the window (for clarity in this section, we drop the frame subscript $t$), for each fine-resolution pixel $(u,v)$, we retrieve the $\kappa\times\kappa$ neighborhood centered around its corresponding coarse location $(u^{\prime},v^{\prime})$, where $u^{\prime}=u/r$ and $v^{\prime}=v/r$ (using subpixel accuracy for $u^{\prime}$ and $v^{\prime}$). The cross-attention map ($CA$) between the fine resolution pixel $(u,v)$ and the pixels $(u_{r},v_{r})$ in its corresponding coarse neighborhood is computed as:
$CA((u,v),(u_{r},v_{r}))=softmax\Big{(}{\bm{q}}(u,v)\cdot{\bm{k}}(u_{r},v_{r})+m\cdot||(u^{\prime},v^{\prime})-(u_{r},v_{r})||_{1}\Big{)},$ (3)
where ${\bm{q}}(u,v)$ is a query representation of the fine-resolution pixel and ${\bm{k}}(u_{r},v_{r})$ is the key representation of its coarse resolution neighbor. The term added to the above dot product represents a static, non-learned spatial bias inspired by Alibi (Press et al., 2022), which demonstrated significantly better generalization to processing text sequences of increasing length through this term. In our case, the bias penalizes attention scores between increasingly distant query-key pairs in terms of their L1 distance. The hyperparameter $m$ (i.e., the slope of the bias) is specific to each head and is set according to Alibi’s configuration.
To extract features needed as input to the query, key, and value transformations of the cross-attention mechanism, we first process the frame through a convnet backbone to extract a feature map at the coarse resolution $\frac{H}{r}\times\frac{W}{r}$. For the fine-resolution pixels used in queries, we further upsample this map using a convnet decoder to obtain denser features. For both query and key locations, we further concatenate their features with the corresponding track features extracted in the coarse resolution. We compute new dense features for the query locations at fine resolution through multiple cross-attention heads using the maps of Eq. 3. Then we transform them through an MLP and a softmax to predict the $\kappa\times\kappa$ weights used in the convex combination of the coarse resolution tracks to compute the tracks at each fine resolution location. We found that more temporally coherent results are produced when the weights are estimated once for the first frame of the time window, and then the same ones are propagated for the rest of the frames.
@@ 4.1 Implementation Details
Training details. We first pretrain the model with 2D loss and visibility loss for 100k iterations, then train with the full loss for another 100k iterations. All stages are conducted on a machine with 8 A100 GPUs. More details are included in the appendix.
@@ 4.2 Comparison to prior works
Baselines. We evaluate our method against prior optical flow and point tracking methods. Particularly, we perform a close comparison against DOT Le Moing et al. (2024) which is a recent SoTA method designed for dense 2D tracking. We implemented a 3D extension of DOT, named DOT-3D, where we incorporate additional depth map input into its optical flow module and add a new head to output $\log(d_{t}/d_{1})$.
@@
[…]
TAPIR + ZoeDepth 9.0 14.3 79.7 5.2 8.8 81.6 10.7 18.2 78.7 8.3 13.8 80.0
BoostTAPIR + ZoeDepth 9.9 16.3 86.5 5.4 9.2 85.3 11.3 19.0 82.7 8.8 14.8 84.8
[…]
SceneTracker + ZoeDepth - 15.1 - - 5.6 - - 16.3 - - 12.3 -
@@
3D point tracking benchmarks results. To further evaluate the generalizability of our approach on in-the-wild videos, we assess its performance on the TAPVid-3D dataset using depth maps estimated by either UniDepth (Piccinelli et al., 2024) or ZeoDepth (Bhat et al., 2023). The results, summarized in Table 5, show that our method consistently outperforms previous approaches, including SpatialTracker, SceneTracker, and 3D-lifted versions of state-of-the-art 2D tracking methods. Our approach demonstrates improvements across most sub-datasets, as well as in the overall average.
@@ 4.3 Ablation Study
Study on the design of upsampler is reported in Tab. 6(c). We compared our approach against upsampling methods using non-learnable operators, such as bilinear, nearest neighbor, and K-nearest neighbor upsamplers. Additionally, we compared our proposed attention-based upsampler with the CNN-based upsampler from RAFT (Teed & Deng, 2020). Our method noticeably outperforms all of these approaches.
@@ Appendix A Appendix
Implementation details. We use the same backbone as Karaev et al. (2023); Harley et al. (2022), which consists of 6 residual blocks, outputting feature maps with a dimension of 256. Unlike CoTracker, which extracts a single-scale feature map and applies pooling later in the correlation module, we directly generate a pyramid of feature maps at scales 2, 4, and 8. We perform dense pixel tracking at a resolution of $H/4\times W/4$ (with $r=4$). The transformer network is composed of 6 spatial and temporal attention blocks, utilizing 8 attention heads and 384 hidden channels. For anchor tracks, we set $9\times 12$ during training and increase to $36\times 48$ during evaluation. In the patch-wise dense local attention, we use a patch size of 6, resulting in $L=36$ tracks per patch. In the high-resolution track upsampler, we use 9 neighbors (with $\kappa=3$) and 2 cross-attention blocks. The number iteration step in the transformer $\Phi$ is set to 6.
We use the AdamW optimizer and the batch size is set to 1 for each GPU. The learning rate is initialized to $10^{-4}$ and scheduled by a linear one cycle (Smith & Topin, 2019). During training, to save the GPU memory consumption, we sample 256 sparse ground-truth trajectories and randomly choose a patch of size $72\times 96$ from the dense 3D trajectory map as supervision. The input video is resized to $384\times 512$ in both training and testing. For in-the-wild video, we leverage ZoeDepth (Bhat et al., 2023) and UniDepth (Piccinelli et al., 2024) to obtain video depth.

### v3

@@ Abstract
Figure 1: DELTA is a dense 3D tracking approach that (a) tracks every pixel from a monocular video, (b) provides consistent trajectories in 3D space, and (c) achieves state-of-the-art accuracy on 3D tracking benchmarks while being significantly faster than previous methods in the dense setting.
@@ 2 Related Work
Optical Flow estimates motion by providing dense pixel-wise correspondences between consecutive frames. Early variational approaches (Mémin & Pérez, 1998; Horn & Schunck, 1980; Brox et al., 2004) struggled with robustness in complex scenes with rapid motion, occlusions, and large displacements. The introduction of CNN-based methods (Ilg et al., 2017; Ranjan & Black, 2017; Xu et al., 2017; Sun et al., 2018) improved motion estimation between adjacent frames. RAFT (Teed & Deng, 2020) marked a breakthrough by leveraging 4D correlation volumes for all pairs of pixels. Follow-up works advanced this by incorporating transformers for tokenizing 4D correlation volumes (Huang et al., 2022), adopting global motion feature aggregation to improve prediction in occluded regions (Jiang et al., 2021), and framing optical flow as a matching problem with correlation softmax operations (Xu et al., 2022). While some efforts propose to apply optical flow to long-term sequences with multi-frame optical flow (Teed & Deng, 2020; Godet et al., 2021; Shi et al., 2023) or integration of point-tracking techniques (Le Moing et al., 2024; Cho et al., 2024a), they often suffer from drifting and occlusion challenges, limiting their reliability for long-term tracking.
@@
We are the first feed-forward approach that performs dense 3D tracking efficiently from a long-term video.  Table 1 provides a brief comparison of existing methods alongside our approach.
@@ 3.1 Preliminary: point tracking
Limitations in dense tracking settings. By default, CoTracker is trained and tested in sparse tracking settings, where the number of tracks $N$ is kept low ($<10^{3}$). In dense settings, where the total number of tracks is $N=H\times W$ (the order of $10^{5}$-$10^{6}$), the spatial attention becomes a bottleneck due to limitations w.r.t. its computational cost and spatial granularity explained below.
Computational complexity. As shown in Fig. 3, applying spatial self-attention across all tokens in a frame (①) results in a computational cost of $TN^{2}$ making it impractical for dense tracking. To reduce complexity, CoTracker introduces virtual track tokens which are conceptually similar to learnable tokens introduced by DETR (Carion et al., 2020). It performs cross-attention back-and-forth between trajectory tokens and a small number of virtual tracks, and self-attention is only applied within the virtual tracks (②). This reduces the computational cost to $2TKN+TK^{2}+KT^{2}$, assuming that the number of virtual tracks $K\ll N$. However, this reduction is still not enough for end-to-end tracking of every pixel in a high-resolution video. Although in practice, pixels could be divided into disjoint groups and perform attention within each group separately, this strategy is sub-optimal due to the interaction between tokens from different groups are ignored.
@@
Method overview. To further reduce computation cost without sacrificing accuracy, we adopt a strategy commonly used in optical flow (Teed & Deng, 2020): performing complex computations at the reduced spatial resolution, followed by lighter layers for upsampling. As shown in Fig. 2, we first run dense tracking on 1/$r^{2}$ of the original resolution, reducing the computation cost by 1/$r^{2}$. Then the reduced-resolution tracks are upsampled to full spatial resolution. In the following section, we use the notation $N=(H\times\ W)/r^{2}$ to denote the number of dense tracks at reduced resolution.
In Sec 3.2, we first discuss our design for reduced-resolution tracking with a new spatial attention architecture, which maintains linear complexity w.r.t. number of tracks while providing finer spatial granularity compared to CoTracker. More importantly, the new architecture can be learned end-to-end for pixel-wise dense tracking without test-train resolution discrepancies. Next in Sec 3.3 we introduce a new transformer-based upsampler that effectively predicts high-res tracking. Finally in Sec 3.4, we delve into details of depth representation that is crucial for accurate tracking in 3D.
@@ 3.2 Joint global and local spatial attention for efficient dense tracking
We design our global-local spatial attention mechanism based on three key criteria: (1) efficiency for both training and testing in dense settings, (2) ability to capture global motion across the image, and (3) ability to capture fine-grained motion details in the neighboring region for each track.
Global attention with sparse anchor tracks. The spatial attention architecture from previous works, i.e. CoTracker (Karaev et al., 2023), which has linear complexity w.r.t. the number of tracks, can perform inference at reduced resolution. However, training remains computationally expensive, exceeding the 80GB GPU memory capacity even when tracking videos with a $96\times 128$ reduced resolution. To improve training efficiency, we employ the following two strategies.
First, a patchwise training strategy is employed to reduce computation in training. At each iteration, we randomly crop small patches of size $h^{\prime}\times w^{\prime}=N^{\prime}$ from the reduced resolution image, then perform dense tracking and obtain supervision within these patches. One issue of patchwise training is that it only computes spatial attention within a local patch without considering the rest region of the first frame. To mitigate this limitation, we augment the patch by introducing a sparse set of $M$ anchor tracks ($M\approx 10^{2}\ll N$), with starting positions uniformly sampled across the first frame.
The second strategy involves computing virtual tracks by cross-attending only to the anchor tracks, instead of attending to all tracks as in CoTracker. As illustrated in ③ of Fig. 3, this approach reduces the cross-attention cost for computing virtual tracks from $TKN^{\prime}$ to $TKM$. Consequently, the total cost for the new global attention becomes $TK(N^{\prime}+2M)+TK^{2}+KT^{2}$, approximately halving the original cost of $2TKN^{\prime}+TK^{2}+KT^{2}$, assuming $T,K,M\ll N^{\prime}$. As shown in Table 6(b), this reduction in computation has minimal impact on tracking accuracy. An additional advantage of learning virtual tracks from anchor tracks is that the same set of anchor tracks is used during both patchwise training and testing on the full image, eliminating any train-test resolution mismatch.
Dense local attention. To capture fine-grained representations of local relations among dense tracks, we apply self-attention within very small spatial patches containing $L$ pixels, prior to cross-attending to the virtual tracks. Since self-attention is applied only within these small patches, this approach adds a marginal complexity of $O(TNL)$ during inference, where $L\ll N$ is the number of tracks per patch. Experiments (Table 6(b)) show that incorporating dense local attention significantly improves dense tracking accuracy.
In summary, our joint Global-Local spatial attention has roughly the same computational cost as CoTracker’s global-only attention but (i) captures both global motion and fined-grained spatial relations, (ii) enables end-to-end training, and (iii) achieves significantly better performance in dense tracking (see Table 2, Table 4 and the qualitative results in our supplementary).
@@ 3.3 High-resolution track upsampler
Given the dense tracks extracted at a reduced spatial resolution of $\frac{H}{r}\times\frac{W}{r}$, the next step is to upsample them to the full resolution $H\times W$. In the context of optical flow, a common upsampling approach expresses the flow for each fine-resolution pixel as a convex combination of its nearest neighbor flows estimated in a coarse resolution (Teed & Deng, 2020). The weights for the combination are learned via a convnet. In contrast, we propose an attention-based upsampling mechanism that more effectively captures the correlation between each fine-resolution pixel and its neighbors at the coarse resolution. We demonstrate the efficacy in our experiments (see Figure 4 and Table 6(c)).
@@
Starting from the first input frame of the window (for clarity, we omit the frame subscript $t$ in this section), the frame is initially processed through a lightweight convolutional backbone to extract a coarse resolution feature map, $\mathcal{F}_{coarse}\in\mathbb{R}^{\frac{H}{r}\times\frac{W}{r}\times D}$, where $D$ is the channel dimension. This coarse feature map is then upsampled using a convolutional decoder to produce a fine-resolution feature map, $\mathcal{F}_{fine}\in\mathbb{R}^{H\times W\times D}$. Each fine-resolution pixel $(u,v)$, with the feature vector $\mathcal{F}_{fine}^{(u,v)}\in\mathbb{R}^{1\times D}$, cross-attends to a $\kappa\times\kappa$ neighborhood centered on its corresponding coarse location $(u^{\prime},v^{\prime})$ in the coarse resolution, where $u^{\prime}=u/r$ and $v^{\prime}=v/r$, using subpixel accuracy for $u^{\prime}$ and $v^{\prime}$. The neighboring coarse features is defined as the set $\{\mathcal{F}_{\text{coarse}}^{(u^{\prime}_{j},v^{\prime}_{j})}\}_{j=1}^{\kappa\times\kappa}\in\mathbb{R}^{(\kappa\times\kappa)\times D}$. Specifically, the fine-resolution feature map is extracted through a cross-attention operation:
$\vspace{-1mm}\mathcal{F}_{fine}^{(u,v)}={\bm{a}}(u,v)\cdot{\bm{v}}({\{\mathcal{F}_{coarse}^{(u^{\prime}_{j},v^{\prime}_{j})}\}_{j=1}^{\kappa\times\kappa})}$ (3)
where the cross-attention scores ${\bm{a}}(u,v)$ are computed as:
$\vspace{-1mm}{\bm{a}}(u,v)=softmax\Big{(}{\bm{q}}(\mathcal{F}_{fine}^{(u,v)})\cdot{\bm{k}}({\{\mathcal{F}_{coarse}^{(u^{\prime}_{j},v^{\prime}_{j})}\}_{j=1}^{\kappa\times\kappa})}+m\cdot||(u^{\prime},v^{\prime})-\{u^{\prime}_{j},v^{\prime}_{j}\}_{j=1}^{\kappa\times\kappa}||_{1}\Big{)}$ (4)
and ${\bm{q}}(\cdot)$, ${\bm{k}}(\cdot),{\bm{v}}(\cdot)$ are linear transformations for the queries, keys, and values respectively. The term added to the above dot product represents a static, non-learned spatial bias inspired by Alibi  (Press et al., 2022). In our case, we bias the query-key attention scores between pixels with a penalty proportional to their distance between their positions (L1 distance in our implementation). We refine the fine-resolution feature map by applying a series of $\tau$ multi-head, Alibi-modified cross-attention blocks, as described in Eq. 3. Finally, we use a MLP to predict the weight map $\mathcal{W}=MLP(\mathcal{F}_{fine})$, where $\mathcal{W}\in\mathbb{R}^{H\times W\times(\kappa\times\kappa)}$. This allows us to compute the high-resolution tracking by taking a weighted average of the coarse-resolution tracks using the predicted weight map. We found that more temporally consistent results are produced when the weights are estimated once for the first frame of the time window, and then the same weights are reused for the rest of the frames.
@@ 4.1 Implementation Details
Training details. Training details are included in the appendix.
@@ 4.2 Comparison to prior works
Baselines. We evaluate our method against prior optical flow and point tracking methods. Particularly, we perform a close comparison against DOT (Le Moing et al., 2024), a recent SoTA method designed for dense 2D tracking. We implemented a 3D extension of DOT, named DOT-3D, where we incorporate depth map input into its optical flow module and add a head to output $\log(d_{t}/d_{1})$.
@@
3D point tracking benchmarks results. To further evaluate the generalizability of our approach on in-the-wild videos, we assess its performance on the TAPVid-3D dataset using depth maps estimated by either UniDepth (Piccinelli et al., 2024) or ZoeDepth (Bhat et al., 2023). The results, summarized in Table 5, show that our method consistently outperforms previous approaches, including SpatialTracker, SceneTracker, and 3D-lifted versions of state-of-the-art 2D tracking methods. Our approach demonstrates improvements across most sub-datasets, as well as in the overall average.
@@ 4.3 Ablation Study
Study on the design of upsampler is reported in Tab. 6(c). We compared our approach against upsampling methods using non-learnable operators (bilinear, nearest neighbor, and 3D K-nearest neighbor) and the CNN-based upsampler from RAFT (Teed & Deng, 2020). Our method noticeably outperforms all of these approaches.
@@ 5 Conclusion
Acknowledgements
Evangelos Kalogerakis has received funding from the European Research Council (ERC) under the Horizon research and innovation programme (Grant agreement No. 101124742).
@@ Appendix A Appendix
Methods Aria DriveTrack PStudio Average
[…]
TAPIR + ZoeDepth 9.0 14.3 79.7 5.2 8.8 81.6 10.7 18.2 78.7 8.3 13.8 80.0
BootsTAPIR + ZoeDepth 9.9 16.3 86.5 5.4 9.2 85.3 11.3 19.0 82.7 8.8 14.8 84.8
TAPTR + ZoeDepth 9.1 15.3 87.8 7.4 12.4 84.8 10.0 17.8 84.3 8.8 15.2 85.6
LocoTrack + ZoeDepth 8.9 15.1 83.5 7.5 12.3 82.8 9.7 17.1 80.1 8.7 14.8 82.1
[…]
SceneTracker + ZoeDepth - 15.1 - - 5.6 - - 16.3 - - 12.3 -
[…]
Table 7: Additional 3D tracking results on the TAP-Vid3D Benchmark. We report the 3D average jaccard (AJ), average 3D position accuracy (APD3D), and occlusion accuracy (OA) across datasets Aria, DriveTrack, and PStudio using ZoeDepth for depth estimation. denotes using depth to lift 2D tracks to 3D tracks.
@@
A.1 Implementation details
Implementation details. We use the same backbone as Karaev et al. (2023); Harley et al. (2022), which consists of 6 residual blocks, outputting feature maps with a dimension of 256. Unlike CoTracker, which extracts a single-scale feature map and applies pooling later in the correlation module, we directly generate a pyramid of feature maps at scales 2, 4, and 8. We perform dense pixel tracking at a resolution of $H/4\times W/4$ (with $r=4$). The transformer network $\Phi$ is composed of 6 spatial and temporal attention blocks, utilizing 8 attention heads and 384 hidden channels. The number iteration step is set to 6. The number of anchor tracks is set to $9\times 12$ during training. In the patch-wise dense local attention, we use a patch size of 6, resulting in $L=36$ tracks per patch. In the high-resolution track upsampler, we use 9 neighbors (with $\kappa=3$) and 2 cross-attention blocks.
Training details. We first pretrain the model with 2D loss and visibility loss for 100k iterations, then train with the full loss for another 100k iterations. All stages are conducted on a machine with 8 A100 GPUs. We use the AdamW optimizer and the batch size is set to 1 for each GPU. The learning rate is initialized to $10^{-4}$ and scheduled by a linear one cycle (Smith & Topin, 2019). During training, to save the GPU memory consumption, we randomly sample a patch of size $N^{\prime}=30\times 40$ from the dense 3D trajectory map as supervision. The input video is resized to $384\times 512$ in both training and testing. For in-the-wild video, we leverage ZoeDepth (Bhat et al., 2023) and UniDepth (Piccinelli et al., 2024) to obtain video depth.
A.2 3D Point Tracking
We additionally report the performance on the TAPVid-3D dataset using depth maps estimated by ZoeDepth (Bhat et al., 2023) in Table 7.
A.3 2D Point Tracking